Paul has relocated and now lives within the 5 kilometre radius. He orders a pizza from the pizza parlour. The probability that it takes between 43 and 49 minutes for him to receive his pizza



1
Expert's answer
2021-10-12T02:23:05-0400

"Given\:that\:\mu =45\:\:and\:\sigma =8"


"P\left(43<x<49\right)=P\left(\frac{43-45}{8}<\frac{\lambda -\mu }{\sigma }<\frac{49-45}{8}\right)"


"=P\left(-0.25<z<0.5\right)"


"=P\left(z<0.5\right)-P\left(z<-0.25\right)"


"=0.6915-0.4013""(From" "z-table)"


"P\left(43<x<49\right)=0.2902"